OpenStudy (anonymous):

How do i find the period of this function? y=cos 1/3x

OpenStudy (anonymous):

period= 2pi/b

OpenStudy (anonymous):

where is b?

OpenStudy (anonymous):

2pi/1/3 so yall gotta multiply 2pi by tha reciprocal its 6pi so yall got 1/3 of the curve in 1 interval

OpenStudy (anonymous):

b=1/3

OpenStudy (anonymous):

18.84 correct ?

OpenStudy (anonymous):

for the period? no.

OpenStudy (anonymous):

6pi

OpenStudy (anonymous):

2∏÷1/3

OpenStudy (anonymous):

therefore multiply 2∏ by the reciprocal of 1/3. 2∏×3

OpenStudy (anonymous):

yeah that's equal 6pi <---

OpenStudy (anonymous):

but how to graph ?

OpenStudy (anonymous):

yes thus the persiod is 6pi

OpenStudy (anonymous):

wolfram it

OpenStudy (anonymous):

well i need to know it for the test :|

OpenStudy (anonymous):

how to graph y=cos1/3x guys ?

OpenStudy (anonymous):

use wolfram.com

OpenStudy (anonymous):

period=2pi/coefficient in front of x. amplitude=whatever trig function is multiplied by.

OpenStudy (kropot72):

\[\cos (1/3)x=\cos (2\pi ft)\] Where \[2\pi f\]is the angular frequency \[(1/3)x=2\pi ft\] \[f=x/(6\pi t)\] Period = T = 1/f \[T=(6\pi t)/x\]

OpenStudy (anonymous):

thats just a complicated way of explaining that and it uses a physics formula that isnt needed for trig

OpenStudy (anonymous):

do you know the easier way to draw this graph ?

OpenStudy (kropot72):

What is your answer for the period Redlinl?

OpenStudy (anonymous):

2pi/coefficient in front of x

OpenStudy (anonymous):

graph it pleaseeeeeeee

OpenStudy (anonymous):



OpenStudy (anonymous):

lol =.=